OpenStudy (anonymous):

For what values of r does the function y= e^rx satisfy the differential equation 2y"+y'-y=0?

OpenStudy (bahrom7893):

do the same thing..

OpenStudy (anonymous):

If r1 and r2 are the values of r that you found in part (a), show that every member of the family of functions =ae^r1x+be^r2x is also a function

OpenStudy (anonymous):

Like I said earlier, my deriving processing is not up to speed yet, Im just trying to get a few questions out so I can learn from understanding the process first

OpenStudy (bahrom7893):

y = e^(rx) y' = re^(rx) y'' = r^2e^(rx)

OpenStudy (bahrom7893):

2y'' + y' - y = 0 2r^2 e^(rx) + r e^(rx) + e^(rx) =0

OpenStudy (bahrom7893):

wait sorry that should be a minus

OpenStudy (bahrom7893):

2y'' + y' - y = 0 2r^2 e^(rx) + r e^(rx) - e^(rx) =0

OpenStudy (bahrom7893):

Divide everything by e^(rx) 2r^2 + r - 1 = 0

OpenStudy (bahrom7893):

r = -1, r = 1/2

OpenStudy (anonymous):

your good haha, mathematics major?

OpenStudy (bahrom7893):

yea

OpenStudy (bahrom7893):

now for the next part: now for the next part: y=ae^r1x+be^r2x y = ae^(-x) + be^(x/2)

OpenStudy (bahrom7893):

y' = -ae^(-x) + (b/2)e^(x/2) y'' = ae^(-x) + (b/4)e^(x/2)

OpenStudy (bahrom7893):

hold up let me write this out on paper, im getting confused..

OpenStudy (anonymous):

so you're going to hate me for this comment but from 2r^2 + r - 1 = 0, did you use the quadratic or did you factor?

OpenStudy (bahrom7893):

I used wolframalpha haha.. and I always use quadratics, I can't factor lol never learned that properly so I don't hate you.. I used http://www.wolframalpha.com/input/?i=Solve%282r%5E2+%2B+r+-+1+%3D+0%29

OpenStudy (anonymous):

heh yess wolfram is a Godsend. Im an atmospheric science major

OpenStudy (anonymous):

still working on the second part?>

OpenStudy (bahrom7893):

Here you go, I get confused if I type, i have to see it on paper..

OpenStudy (anonymous):

thank you

OpenStudy (bahrom7893):

np